You have been managing a $5 million portfolio that has a beta of 1.25 and a required rate of return of 8.875%. The current risk-free rate is 2%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 1.45, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Currently:

8.875=2+1.25*(market rate-2)

(8.875-2)=1.25(market rate-2)

market rate=(8.875-2)/1.25+2

=7.5%

Hence required rate for $500,000=2+1.45*(7.5-2)=9.975%

Portfolio return=Respective return*Respective weight

=(5,000,000/5,500,000*8.875)+(500,000/5,500,000*9.975)

=8.98%(Approx).